OpenStudy (anonymous):

If 5x2 + y4 = −9 then evaluate the second derivative of y with respect to x when x = 2 and y = 1

OpenStudy (loser66):

The easiest way is isolate y

OpenStudy (loser66):

\(y = (-5x^2-9)^{1/4}\) then take derivative twice, then apply x=2 in

OpenStudy (loser66):

but if they give you x =2 and y =1, that means they want you to find implicit derivative. ok, just do it

OpenStudy (anonymous):

So where would I apply the y? They do want me to find the implicit derivative

OpenStudy (loser66):

do it, just do as usual. derivative both sides, right hand side =0, right? \(10x+4y^3y'=0\)

OpenStudy (loser66):

hence \(y'= \dfrac{-10x}{4y^3}\)

OpenStudy (loser66):

let it there, back to the first one, take second derivative

OpenStudy (loser66):

\(10 +12y^2y'+4y^3y"=0\)

OpenStudy (loser66):

isolate y", you have x, y, y', replace all to get the answer

OpenStudy (loser66):

Next time, if I find out you disappear when I am trying to help you figure out the problem, I will delete my guidance. Please, respect the helpers. We are here for free.